OpenStudy (anonymous):

HELP please no one else will help me 1 3/4% is the tax. $43210 annually: Find the local tax deducted

OpenStudy (anonymous):

1-3/4%=.0175

OpenStudy (anonymous):

can you retype the tax? I don't know if it's correctly written.

OpenStudy (anonymous):

Right, The question is not clearly written.. No wonder no one is helping.

OpenStudy (anonymous):

what @benzilla said \[1\tfrac{3}{4}\%=.0175\]

OpenStudy (anonymous):

yeah well its not my fault thats how my teacher gave it too all of us

OpenStudy (anonymous):

your job is to take a calculator and compute \[.0175\times 43210 \] if i am reading the question correctly

OpenStudy (anonymous):

can you help with this too ? Find the amount of social security deducted from the check: $835 biweekly

OpenStudy (anonymous):

i get $756.175 which you should probably round to $756.18 since it is money

OpenStudy (anonymous):

you mean yearly amount deducted? there are 52 week in a year, so take \[835\times 26\]

OpenStudy (anonymous):

no its BiWEEKLY, ig it means every two weeks

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

$756.17